# bibliography.bib

### v1

@@ 3.2 Task
In this paper, we aim to extend PFNs to causal inference. Here, the main challenge is that the object of interest is an interventional [^2] Causal literature often distinguishes between interventional and counterfactual distributions. This is not relevant for the methods of our paper, and we thus use interventional distribution as an umbrella term that incorporates both. distribution $\mathbb{P}_{\mathrm{int}}$, yet we only observe data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}$ from a potentially different observational distribution [Pearl.2009].
@@
As an illustrative example, we consider a standard causal inference setting where the data comprise $(X,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$, where $X$ are patient covariates, $A$ is a treatment and $Y$ an outcome of interest [vanderLaan.2006, Shalit.2017]. For example, in medicine, $X$ may contain treatment history or demographic attributes $A$ may be a medical treatment, and $Y$ a health outcome. Following the potential outcome framework [Rubin.1974], let $Y(a)$ denote the outcome that would realize under the intervention $\mathrm{do}(A=a)$. The interventional distribution is thus over $(X,A,Y(1)-Y(0))\sim\mathbb{P}_{\mathrm{int}}$, and a common target functional is the conditional average treatment effect (CATE) $Q(x)\;=\;\mathbb{E}\bigl{[}Y(1)-Y(0)\mid X=x\bigr{]}$ [Wager.2018]. The CATE quantifies the expected benefit of providing treatment given the patient’s covariates.
Identifiability. To estimate CATE from observational data, we need to impose identifiability assumptions, which link the observational and the interventional distributions and allow us to express $Q$ as a functional of the observational distribution [Rosenbaum.1983]. These are (i) consistency: $Y(A)=Y$, (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)>0$, and (iii) Unconfoundedness: $Y(1),Y(0)\perp\!\!\!\perp A\mid X$ in $\mathbb{P}_{\mathrm{int}}$.
Instead of Assumption (iii), weaker assumptions are often made (e.g., assuming an instrumental variable while allowing for unconfoundedness). We unify these different assumptions and causal queries in the following.
@@
We define a causal inference setting as a tuple
$\mathcal{C}\;=\;\bigl{(}Z,\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}},\,Q\bigr{)},$ (6)
where $Z$ collects the observed variables (and contains at least $A$ and $Y$); $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}$ are paired observational/interventional distributions over $Z$ that correspond to an intervention on $A$; and $Q(\mathbb{P}_{\mathrm{int}})$ is a causal query that is identifiable, i.e. there exists a measurable functional $\bar{Q}$ such that
$Q(\mathbb{P}_{\mathrm{int}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}})\quad\text{for all }(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}.$ (7)
@@ 3.2.1 Running examples
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy (i) consistency: $Y=Y(A,M)$ and $M=M(A)$; (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=a\mid X=x)>0$ and $\mathbb{P}_{\mathrm{obs}}(M=m\mid A=a,X=x)>0$; and (iii) front-door criterion $M(a)\;\perp\!\!\!\perp\;A\mid X=x,\quad\text{and}\quad Y(a^{\prime},m)\;\perp\!\!\!\perp\;M\mid A=a^{\prime},\,X=x$. Under these assumptions, the CATE is identified via
$\widetilde{Q}(\mathbb{P}_{\mathrm{obs}})(x)$ $\;=\;\sum_{m\in\mathcal{M}}\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{[}Y\mid A=1,M=m,X=x\bigr{]}\,\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)\,\mathbb{P}_{\mathrm{obs}}(M=m\mid A=1,X=x)$ (9)
$-\sum_{m\in\mathcal{M}}\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{[}Y\mid A=0,M=m,X=x\bigr{]}\,\mathbb{P}_{\mathrm{obs}}(A=0\mid X=x)\,\mathbb{P}_{\mathrm{obs}}(M=m\mid A=0,X=x),$ (10)
which is the conditional version of Pearl’s front-door adjustment [Pearl.2009], thereby proving that $Q(\mathbb{P}_{\mathrm{int}})=\widetilde{Q}(\mathbb{P}_{\mathrm{obs}})$ for all admissible pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$.
@@
$Q(\mathbb{P}_{\mathrm{int}})(x,a)\;=\;\mathbb{E}\!\bigl{[}Y(a)\mid X=x\bigr{]},$ (11)
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy the following conditions [Newey.2003]: (i): Additive structural equation: $Y=f(X,A)+g(X,U)$, with an unknown functions $f$ and $g$ and unobserved confounder $U$, implying that $Y$ does not directly depend on $Z$; (ii) Independence: $U\perp\!\!\!\perp Z\mid X$; (iii) Relevance: $\mathbb{P}_{\mathrm{obs}}(A\mid X=x,Z=z)>0$ is non-constant in $z$; and (iv) Completeness: For every measurable $g$, if $\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[g(A,x)\mid X=x,Z=z]=0$ for all $z$, then $g(A,x)=0$ almost surely in $\mathbb{P}_{\mathrm{obs}}$. Then, the CAPO can be written as $Q(\mathbb{P}_{\mathrm{int}})(x,a)=f(x,a)+\mathbb{E}_{\mathbb{P}_{\mathrm{int}}}[g(x,U)\mid X=x]$ and is identified via the integral equation
$\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Y\mid X=x,Z=z]$ $=\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[f(X,A)\mid X=x,Z=z]+\mathbb{E}_{\mathbb{P}_{\mathrm{int}}}[g(X,U)\mid X=x]$ (12)
$=\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Q(\mathbb{P}_{\mathrm{int}})(X,A)\mid X=x,Z=z],$ (13)
which has a unique solution $\bar{Q}(\mathbb{P}_{\mathrm{obs}})(x,a)$ due to the completeness condition (iv) [Newey.2003, Hartford.2017].
Research question: PFNs have shown to be an effective way to construct tabular foundation models. However, a causal inference setting $\mathcal{C}$ comes with additional challenges, such as the distinction between observational and interventional distribution ans well as identifiability assumptions. Our main research question is:
@@
Below, we introduce our CausalML consisting of (1) appropriate priors (see Section 4) and (2) a training algorithm for the underlying PNFS (see Section 5).
@@ 4 CausalFM: Priors
The first step to derive our CausalML (and thus train a PFN) is to construct a valid prior distribution $\Pi$ that allows for Bayesian inference and approximating the predictive posterior distribution as in Eq. (5). In this section, we discuss how to construct such priors for a causal inference setting.
4.1 Motivation for SCM-based priors
@@
We motivate our approach to constructing priors with a normally distributed IV setting. That is, we consider Example 3 from above (ignoring $X$) where $(Z,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$ is a multivariate normal distribution with mean zero. A straightforward way would be to put a prior on the parameters (covariance matrix) of this distribution, thus inducing a prior $\Pi$ on the observational distribution $\mathbb{P}_{\mathrm{obs}}$. If the posterior $\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n})\to\mathbb{P}_{\mathrm{obs}}^{\ast}$ converges to the ground-truth observational distribution $\mathbb{P}_{\mathrm{obs}}^{\ast}$ (i.e., satisfying a Bernstein-von-Mises theorem), we can obtain a consistent Bayesian estimator of our causal query via $\bar{Q}(\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n}))$. Accordingly, we could train a PFN $q_{\theta}(Y\mid\mathcal{D}_{n},a)$ with the loss in Eq. (5) and estimate the CAPO via $\bar{Q}(q_{\theta}(Y\mid\mathcal{D}_{n},a))$.

### v2

@@ 3.2 Task: causal inference
In this paper, we aim to extend PFNs to causal inference. Here, the main challenge is that the object of interest is an interventional [^2] Causal literature often distinguishes between interventional and counterfactual distributions. This is not relevant for the methods of our paper, and we thus use interventional distribution as an umbrella term. distribution $\mathbb{P}_{\mathrm{int}}$, yet we only observe data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}$ from a potentially different observational distribution (Pearl, 2009).
@@
As an illustrative example, we consider a standard causal inference setting, called backdoor adjustment, where the data comprise $(X,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$, where $X$ are patient covariates, $A$ is a treatment, and $Y$ is an outcome of interest (van der Laan & Rubin, 2006). For example, in medicine, $X$ may contain treatment history or demographic attributes, $A$ may be a medical treatment, and $Y$ a health outcome. Following the potential outcome framework (Rubin, 1974), let $Y(a)$ denote the outcome that would be realized under the treatment $A=a$. The interventional distribution is thus over $(X,A,Y(1)-Y(0))\sim\mathbb{P}_{\mathrm{int}}$, and a common target functional is the conditional average treatment effect (CATE) $Q(x)\;=\;\mathbb{E}\bigl[Y(1)-Y(0)\mid X=x\bigr]$ (Wager & Athey, 2018). The CATE quantifies the expected benefit of providing treatment given the patient’s covariates.
Identifiability. To estimate CATE from observational data, we need to impose identifiability assumptions, which link the observational and the interventional distributions and allow us to express $Q$ as a functional of the observational distribution (Rosenbaum & Rubin, 1983). These are (i) consistency: $Y(A)=Y$, (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)>0$, and (iii) Unconfoundedness: $Y(1),Y(0)\perp\!\!\!\perp A\mid X$ in $\mathbb{P}_{\mathrm{int}}$.
@@
We define a causal inference setting is a tuple $\mathcal{C}\;=\;\bigl(O,\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}},\,Q\bigr)$, where $O$ collects the observed variables (and contains at least $A$ and $Y$); $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}$ are paired observational/interventional distributions over $O$ that correspond to an intervention on $A$; and $Q(\mathbb{P}_{\mathrm{int}})$ is a causal query that is identifiable, i.e. there exists a measurable functional $\bar{Q}$ such that $Q(\mathbb{P}_{\mathrm{int}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}})$ for all $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}$.
@@ 3.2.1 Running examples
Figure 1: C-DAGs compatible with the three example causal inference settings. Yellow variables are observed, blue variables are unobserved, and red variables are clusters of variables.
@@
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy (i) consistency: $Y=Y(A,M)$ and $M=M(A)$; (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=a\mid X=x)>0$ and $\mathbb{P}_{\mathrm{obs}}(M=m\mid A=a,X=x)>0$; and (iii) front-door criterion $M(a)\;\perp\!\!\!\perp\;A\mid X=x,\quad\text{and}\quad Y(a^{\prime},m)\;\perp\!\!\!\perp\;M\mid A=a^{\prime},\,X=x$. Under these assumptions, the CATE is identified and $\widetilde{Q}$ is given via the conditional version of Pearl’s front-door adjustment formula (Pearl, 2009).
@@
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy the following conditions (Newey & Powell, 2003): (i): Additive structural equation: $Y=f(X,A)+g(X,U)$, with (unknown) functions $f$ and $g$ and unobserved confounder $U$, implying that $Y$ does not directly depend on $Z$; (ii) Independence: $U\perp\!\!\!\perp Z\mid X$; (iii) Relevance: $\mathbb{P}_{\mathrm{obs}}(A\mid X=x,Z=z)>0$ is non-constant in $z$; and (iv) Completeness: For every measurable $g$, if $\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[f(x,A)\mid X=x,Z=z]=0$ for all $z$, then $f(x,A)=0$ almost surely in $\mathbb{P}_{\mathrm{obs}}$. Then, the CATE can be shown to be identified via an integral equation (Newey & Powell, 2003; Hartford et al., 2017).
Research question: PFNs have shown to be an effective way to construct tabular foundation models. However, a causal inference setting $\mathcal{C}$ comes with additional challenges, such as the distinction between observational and interventional distribution as well as identifiability assumptions.
@@
In the following, we introduce CausalFM consisting of (i) appropriate prior distributions that allow for approximating interventional predictive posterior distributions as in Eq. (2)(Section 4) and (ii) a training algorithm for the underlying PNFS (see Section 5).
@@ 4 CausalFM: Priors
In this section, we construct prior distributions for CausalFM which are based upon identifiable structural causal models (SCMs). We motivate and formalize our approach (Sec. 4.1, provide necessary criteria for valid causal inference (Sec. 4.2), and finally provide a method for constructing such priors in practice (Sec. 4.3). We also provide a complete toy example in Appendix B.
4.1 Introducing SCM-based priors
@@
Naïve approach. A naïve approach for causal inference would construct a prior $\Pi$ directly for the observational distribution $\mathbb{P}_{\mathrm{obs}}$. If the posterior $\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n})\to\mathbb{P}_{\mathrm{obs}}^{\ast}$ converges to the ground-truth observational distribution $\mathbb{P}_{\mathrm{obs}}^{\ast}$ (i.e., satisfying a Bernstein-von-Mises theorem), we can obtain a consistent Bayesian estimator of our causal query via $\bar{Q}(\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n}))$. Accordingly, we could train a PFN $q_{\theta}(Y\mid\mathcal{D}_{n},x)$ with the loss in Eq. (2) and estimate the CAPO via $\bar{Q}(q_{\theta}(Y\mid\mathcal{D}_{n},x))$.